Hence solve the /VP. 12 Use a similar approach to Questions 8 and 9 to solve these IVPs. DX = with y(t) = 2 with y(2) = 1 1 y Q12 f only. please.
Answer #1

(2 - x Given - ② Ø Giren dy = Y(2-x) om die (2) dx Integrating Idy = J 2 hor din - J du 5) : ln (y) = -3 - In (x3 + -O CcinaAlternative form enlys =-2 / - Incantl = lnly) t. ln (2) = - 2 + 1 = ln (ya) = (1 - 2 / 2 ) =) ya = (1 - 2 / 2 7 y=1ec1- 2 )